an airplane flies eastward and accelerate

An airplane flies eastward and accelerates uniformly. At one position along its path it has a velocity of 30.3 m/s. It then flies a further distance of 42.9 km and its velocity is 41.5 m/s. Find the airplanes acceleration and calculate how much time elapses while the airplane covers those 42.9 km Acceleration: Number m/s Elapsed time: Number
Answer #1

Solution)

We know, v^2=u^2 + 2as

41.5^2 =30.3^2 + 2*a*42900

Acceletation, a= 0.00624 m/s^2

=====

Now, v= u+ at

41.5=30.3 + 0.00624*t

Time, t= 1794.87 secs

======